# Qubit encoding for a mixture of localized functions

### v1

@@ III Experiments
Figure 9 shows the observed probability distributions for $n_{q}=4$ data qubits and one ancilla according to probabilistic encoding of $|\psi_{\mathrm{lc}}\rangle\propto|L;a_{0},k_{\mathrm{c}0}\rangle_{n_{q}}+|L;a_{1},k_{\mathrm{c}1}\rangle_{n_{q}}$ with $a_{0}=a_{1}=0.5,k_{\mathrm{c}0}=0,$ and $k_{\mathrm{c}1}=8.$ The ideal distribution is also shown on each panel. It is seen that the observed distributions and the ideal one are in reasonable agreement for all the eight quantum computers. The number of CNOT gates in the transpiled circuit was 23.
@@ IV Conclusions
In summary, we developed a probabilistic encoding technique that generates an LC of LFs with arbitrary complex coefficients. Since the ancillae increase only logarithmically with respect to the number of constituent LFs regardless of the size of data register and QFT is performed only once, this technique achieves efficient encoding. Furthermore, the QARA technique was shown to determinize the encoding technique with controllable errors. We performed the probabilistic encoding technique on commercially available real quantum computers of today and found that the results were in reasonable agreement with the theoretical ones.
Looking beyond the noisy intermediate-scale quantum device (NISQ) era, one of the intriguing applications of our generic scheme will be the encoding of MOs used in quantum chemistry in real space. In particular, we found via resource estimation that the computational time for encoding a localized MO is polynomial in terms of the logarithm of electron number. These insights are encouraging since the encoding of MOs is a crucial part for the state preparation of an initial many-electron wave function. The encoding techniques developed in the present study thus make quantum chemistry in real space more promising on fault-tolerant quantum computers.
@@ C.1 $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$
We first consider $X$ gates on $m_{\mathrm{t}}$ target qubits $(m_{\mathrm{t}}\geq 2)$ controlled by $m_{\mathrm{c}}$ qubits, as depicted on the left-hand side of Fig. 11(a). We adopt the same technique for this $(m_{\mathrm{c}}+m_{\mathrm{t}})$-qubit gate as in the original method [50]. We describe the technique here in order for this paper to be self contained.
We define partial circuits $S_{k}$ for $k=1,\dots,\lceil\log_{2}m_{\mathrm{t}}\rceil$ as shown on the right-hand side of Fig. 11(a). $S_{k}$ for each $k$ is composed only of at most $2^{k-1}$ simultaneously executable CNOT gates. One can easily confirm the equality in Fig. 11(a), that is, $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ can be implemented by arranging the $S_{k}$ circuits in descending order, appending a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate, and arranging the $S_{k}$ circuits in ascending order. Since the $\mathrm{C}^{m_{\mathrm{c}}}X$ gate admits a linear-depth implementation [45, 46], the total depth is given by
@@ C.2 MCM1 gates
where we used the relation $\sigma_{x}Z_{\ell}^{1/2\dagger}\sigma_{x}=e^{-i\theta_{\ell}/2}Z_{\ell}^{1/2}.$ Equation (56) confirms that the equality in Fig. 11(b) holds regardless of the ancillary state. The only many-qubit gates in $\mathcal{C}$ are $\mathrm{C}^{n_{\mathrm{A}}-1}Z(\widetilde{\alpha})$ and $\mathrm{C}^{n_{\mathrm{A}}}X^{\otimes n_{q}}.$ The former can be implemented with a linear depth [45, 46] and the latter can be implemented using the technique shown in Fig. 11(a). The total depth for the MCM1 gates is thus
@@
C.3 Implementation of QFT using the phase gradient method
As pointed out in Ref. [50], the phase gradient method can be employed for efficient implementation of QFT. We provide here the circuit diagram of the efficient implementation for this paper to be self contained.
For an $n$-qubit system, we define phase gates $R_{k}\equiv Z(2\pi/2^{k})$ for $k=2,\dots,n.$ By using $V^{(k)}$ partial circuits defined in Fig. 12(a) as building blocks, we construct the circuit $\mathcal{C}_{\mathrm{PG-QFT}}$ in Fig. 12(b). This circuit implements QFT for the $n$ qubits. One can confirm that, by rearranging the phase gates in the standard implementation [43], $\mathcal{C}_{\mathrm{PG-QFT}}$ is obtained and it thus works as a QFT circuit. Since the depth of $V^{(k)}$ is $\mathcal{O}(\log k)$ for each $k$ thanks to the technique in Figs. 11(a) and (b), we have $\mathrm{depth}(\mathcal{C}_{\mathrm{PG-QFT}})=\mathcal{O}(n\log n).$
Figure 12: (a) Definition of partial circuit $V^{(k)}$ for $k=2,\dots,n.$ (b) Circuit $\mathcal{C}_{\mathrm{PG-QFT}}$ for QFT based on the phase gradient method.
@@ Appendix D Formulation for an LC of complex localized functions via complex coefficients
Let us consider a case where we want to encode an LC of localized complex functions $\{f_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}$ via complex coefficients $\{d_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}.$ By separating the real and imaginary parts, we write $f_{\ell}=f^{(\mathrm{R})}_{\ell}+if^{(\mathrm{I})}_{\ell}$ and $d_{\ell}=|d_{\ell}|e^{i\vartheta_{\ell}}$ for each $\ell.$ We assume that the unitaries $U^{(\ell,\mathrm{R})}$ and $U^{(\ell,\mathrm{I})}$ that generate $f^{(\mathrm{R})}_{\ell}$ and $f^{(\mathrm{R})}_{\ell},$ respectively, at the origin are known.

### v2

@@ III Experiments
Figure 8 shows the observed probability distributions for $n_{q}=4$ data qubits and one ancilla according to the probabilistic encoding of $|\psi_{\mathrm{lc}}\rangle\propto|L;a_{0},k_{\mathrm{c}0}\rangle_{n_{q}}+|L;a_{1},k_{\mathrm{c}1}\rangle_{n_{q}}$ with $a_{0}=a_{1}=0.5,k_{\mathrm{c}0}=0,$ and $k_{\mathrm{c}1}=8.$ The ideal distribution is also shown on each panel. It is seen that the observed distributions and the ideal one are in reasonable agreement for all the eight quantum computers. The number of CNOT gates in the transpiled circuit was 23.
@@ IV Conclusions
In summary, we developed a probabilistic encoding technique that generates an LC of LFs with arbitrary complex coefficients. Since the ancillae increase only logarithmically with respect to the number of constituent LFs regardless of the size of the data register and QFT is performed only once, this technique achieves efficient encoding. Furthermore, the QARA technique was shown to determinize the encoding technique with controllable errors. We performed the probabilistic encoding technique on commercially available real quantum computers and found that the results were in reasonable agreement with the theoretical ones.
Looking beyond the noisy intermediate-scale quantum device (NISQ) era, one of the intriguing applications of our generic scheme will be the encoding of MOs used in quantum chemistry in real space. In particular, we found via resource estimation that the computational time for encoding a localized MO is polynomial in terms of the logarithm of the electron number. These insights are encouraging since encoding MOs is crucial for the state preparation of an initial many-electron wave function. The encoding techniques developed in the present study thus make quantum chemistry in real space more promising on fault-tolerant quantum computers. Also, the techniques will be a powerful tool in the field of financial analyses using quantum computers, where they often want to encode various probability distribution functions in continuous space efficiently.
@@ C.1 $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$
We first consider $X$ gates on $m_{\mathrm{t}}$ target qubits $(m_{\mathrm{t}}\geq 2)$ controlled by $m_{\mathrm{c}}$ qubits, as depicted on the left-hand side of Fig. 10(a). This gate is known as the quantum fan-out gate [55, 56]. We adopt the same technique for this $(m_{\mathrm{c}}+m_{\mathrm{t}})$-qubit gate as in the original method [54]. We describe the technique here in order for this paper to be self contained.
We define partial circuits $S_{k}$ for $k=1,\dots,\lceil\log_{2}m_{\mathrm{t}}\rceil$ as shown on the right-hand side of Fig. 10(a). $S_{k}$ for each $k$ is composed only of at most $2^{k-1}$ simultaneously executable CNOT gates. One can easily confirm the equality in Fig. 10(a), that is, $\mathrm{C}^{m_{\mathrm{c}}}X^{\otimes m_{\mathrm{t}}}$ can be implemented by arranging the $S_{k}$ circuits in descending order, appending a $\mathrm{C}^{m_{\mathrm{c}}}X$ gate, and arranging the $S_{k}$ circuits in ascending order. Since the $\mathrm{C}^{m_{\mathrm{c}}}X$ gate admits a linear-depth implementation [48, 49, 50], the total depth is given by
@@ C.2 MCM1 gates
where we used the relation $\sigma_{x}Z_{\ell}^{1/2\dagger}\sigma_{x}=e^{-i\theta_{\ell}/2}Z_{\ell}^{1/2}.$ Equation (48) confirms that the equality in Fig. 10(b) holds regardless of the ancillary state. The only many-qubit gates in $\mathcal{C}$ are $\mathrm{C}^{n_{\mathrm{A}}-1}Z(\widetilde{\alpha})$ and $\mathrm{C}^{n_{\mathrm{A}}}X^{\otimes n_{q}}.$ The former can be implemented with a linear depth [48, 49, 50] and the latter can be implemented using the technique shown in Fig. 10(a). The total depth for the MCM1 gates is thus
@@ Appendix D Formulation for an LC of complex localized functions via complex coefficients
Let us consider a case where we want to encode an LC of localized complex functions $\{f_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}$ via complex coefficients $\{d_{\ell}\}_{\ell=0}^{n_{\mathrm{loc}}-1}.$ By separating the real and imaginary parts, we write $f_{\ell}=f^{(\mathrm{R})}_{\ell}+if^{(\mathrm{I})}_{\ell}$ and $d_{\ell}=|d_{\ell}|e^{i\vartheta_{\ell}}$ for each $\ell.$ We assume that the unitaries $U^{(\ell,\mathrm{R})}$ and $U^{(\ell,\mathrm{I})}$ that generate $f^{(\mathrm{R})}_{\ell}$ and $f^{(\mathrm{I})}_{\ell},$ respectively, at the origin are known.